A valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.  jegy

diákok:

Ferenczik Judit  +  +  +  ±  +  +  5
Husiová Anikó  −  −  +  −  +  ±
Kiss Andrea Erzsébet  +  +  +  −  +  +  4
Mikó Tamás  ±+  +  ±  −  ±  −  2
Nemoda Dóra  +  +  +  −  +  +  4
Porvázsnyik Bettina  +  +  ±  −  +  +  3
Szabó Lívia Dóra  +  +  +  −  +  +  4